OpenStudy (anonymous):

use a half angle formula to find the exact value of each expression tan 7pi/8

OpenStudy (anonymous):

@aum

OpenStudy (anonymous):

\[\large \sin^2{x}=\frac{ 1-\cos2x }{ 2 }\] \[\large \cos^2{x}=\frac{1+\cos2x}{2}\]

OpenStudy (anonymous):

i got it thanks

OpenStudy (anonymous):

Alright